#opnfv-meeting: BGS daily release readiness synch

Meeting started by frankbrockners at 15:59:54 UTC (full logs).

Meeting summary

    1. Frank Brockners (frankbrockners, 16:00:01)
    2. morgan (morgan_orange, 16:01:32)
    3. Manuel (ManuelR_Sandvine, 16:02:09)
    4. Tim Rozet (trozet, 16:02:14)
    5. pbandzi (pbandzi, 16:02:33)
    6. Heather Kirksey (HKirksey, 16:02:55)
    7. Chris Price (late as usual :) (ChrisPriceAB, 16:04:10)
    8. topics are as usual: check where we are - and reflect latest status on https://wiki.opnfv.org/releases/arno/releasereadiness (frankbrockners, 16:04:47)
    9. Trevor Cooper (trevor_intel, 16:05:45)
    10. special attention needs to go to the "target system state topic" https://wiki.opnfv.org/get_started/get_started_system_state (frankbrockners, 16:05:50)
    11. Let's start with system state (frankbrockners, 16:06:04)
    12. Functest needs some answers - so that they can decide whether certain tests are feasible (frankbrockners, 16:06:31)
    13. several entries on ostack are "not confirmed" - anyone have updates on whether we will have them available? (frankbrockners, 16:07:15)
    14. what's an ostack entry referring to? ah ok (ChrisPriceAB, 16:08:08)
    15. heat used for vIMS (morgan_orange, 16:08:11)
    16. Foreman installs Horizon, but not Heat or Ceilometer (trozet, 16:08:49)
    17. Shouldn't be too bad to enable them if needed, just need time to make changes and test that its working (trozet, 16:09:15)
    18. do we agree that those 2 are both needed, or just Heat? (trozet, 16:09:30)
    19. do we have new requirements coming in days from release? (ChrisPriceAB, 16:10:06)
    20. aside form some consternation about that, the fuel deploy comes with heat (ChrisPriceAB, 16:10:39)
    21. looks like to me ... - IMHO we should stay with what we have for now. If we have sparetime - we can revisit new components. Thoughts? (frankbrockners, 16:10:50)
    22. https://wiki.opnfv.org/vims_functional_testing (morgan_orange, 16:11:08)
    23. ack (ChrisPriceAB, 16:11:09)
    24. Agree. After LF hardware is working, then I can focus on adding Heat for Foreman Installer (trozet, 16:11:23)
    25. I'll update that wiki page today (trozet, 16:11:39)
    26. let's not make it a requirement. If the guys can swing it up (once we are in stable state) without significant delays we might want to. (ChrisPriceAB, 16:11:47)
    27. Agreed. Let's not add mandatory things to the plate the last minute. Per Tim: We'll consider adding Heat - time permitting (so low prio for now). (frankbrockners, 16:13:05)
    28. Cinder/Ceph and Swift are also missing form the components list. Will update that on the wiki. (trozet, 16:13:11)
    29. quick status update / progress made since yesterday (frankbrockners, 16:14:04)
    30. Foreman was still failing to deploy on LF hardware. I found there was a VLAN tag with VID 0 being added to frames coming back to the host (trozet, 16:14:46)
    31. pbandzi helped debug. We are still not sure why that tag is added by the UCS/FI, but we can fix the problem by restarting networking inside of the VM (trozet, 16:15:18)
    32. big thanks to pbandzi for finding the work around. Will update our deploy script to include the workaround (trozet, 16:15:45)
    33. so current status now is trying to deploy again with the fix (trozet, 16:16:28)
    34. radez also has a fix for the ISO which will make it bootable (trozet, 16:16:49)
    35. On the fuel track side they have turned a lot of attention to working with getting the functest activities running against the Arno deploy (ChrisPriceAB, 16:17:43)
    36. http://ircbot.wl.linuxfoundation.org/meetings/opnfv-bgs/2015/opnfv-bgs.2015-04-23-13.59.html - report of the recent team meeting on Fuel related issues (frankbrockners, 16:18:07)
    37. on the deploy side the metal deploy is working on the community lab, networking description is being prepared for porting to the LF lab (ChrisPriceAB, 16:18:26)
    38. next step is to get the NW in place tidy up the deploy and start working in the LF lab. (ChrisPriceAB, 16:18:50)
    39. pending pbandzi getting the fuel NW peices in place in order to keep the whole ship moving (ChrisPriceAB, 16:19:12)
    40. pending pbandzi getting the foreman NW pieces in place in order to keep the whole ship moving (ChrisPriceAB, 16:19:29)
    41. functest VM compilation ongoing for deployment to JH for automated validation (ChrisPriceAB, 16:20:11)
    42. Jonas to contact pbandzi to discuss current states and migration activities (ChrisPriceAB, 16:21:09)
    43. Updates from FuncTest - especially: Do we have a definition of which tests are considered for Arno? (frankbrockners, 16:22:39)
    44. bench rally ok (manual and CI test) - to be tested on E/// POD and full list of scenario to be finalized (morgan_orange, 16:22:52)
    45. tempest rally ok manual but ko for Ci (jenkins job pending => to be analyzed) (morgan_orange, 16:23:27)
    46. ODL pbandzi commited some code tonight, it shall be not far from being OK (morgan_orange, 16:23:55)
    47. vPing no news of the comiter today, open question regarding the Tempest based scenario (morgan_orange, 16:24:27)
    48. Does functest expect to run any other tests than a subset of tempest scenario tests and robot tests retrieved from either openstack or ODL trunk? (frankbrockners, 16:25:04)
    49. vIMS, contributors did not commit stuff yet, they are waiting for a stable env to test their solution (that is fine on their lab) (morgan_orange, 16:25:06)
    50. were not able to upload images, initiate VMs, (morgan_orange, 16:25:29)
    51. jose_lausuch will help to fix issue on E/// POD (prio 1) before moving to LF (morgan_orange, 16:26:10)
    52. I told them to start document in the rst and commit their script (morgan_orange, 16:26:30)
    53. only test for R1 which is neither from ODL or OSTACK trunk is vIMS test (frankbrockners, 16:28:22)
    54. Do we require any specific VNFs for R1 testing? And if so where do they come from? (frankbrockners, 16:28:47)
    55. for R1 only specific VNFS should have been vPing scenario and vIMS (with a VNF IMS + a VNF SIP loader) (morgan_orange, 16:29:45)

Meeting ended at 16:31:41 UTC (full logs).

Action items

  1. (none)

People present (lines said)

  1. frankbrockners (39)
  2. ChrisPriceAB (27)
  3. morgan_orange (22)
  4. trozet (17)
  5. collabot (3)
  6. dneary (2)
  7. trevor_intel (1)
  8. ManuelR_Sandvine (1)
  9. pbandzi (1)
  10. HKirksey (1)

Generated by MeetBot 0.1.4.